ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 158, 2007

Abert, M. and L. Babai Finite groups of uniform logarithmic diameter

193

Babai, L. See M. Abert

Barvinok, A. and A. Samorodnitsky Random weighting, asymptotic counting, and inverse isoperimetry

159

Bayart, F. and L. Quarta Algebras in sets of queer functions  285

Begun, B. and A. del Junco Amenable groups, stationary measures and partitions with independent iterates

41

Bell, M. and W. Marciszewski On scattered Eberlein compact spaces

217

Bentkus, V. On measure concentration for separately Lipschitz functions in product spaces

1

Buckley, S. M. and D. A. Herron Uniform domains and capacity 129
Ershov, M. Finite presentability of SL_{1}(D)  297

Giambruno, A., D. La Mattina and V. M. Petrogradsky Matrix algebras of polynomial codimention growth

367

Haydon, R., A. Moltó and J. Orihuela Spaces of functions with countably many discontinuities

19

Herron, D. A. See S. M. Buckley

Hochman, M. Averaging sequences and abelian rank in amenable groups

119

Huang, W., X. D. Ye and G. Zhang Relative entropy tuples, relative u.p.e. and c.p.e. extensions

249

del Junco, A. See B. Begun

Kerner, D. Erratum for: “Enumeration of singular algebraic curves” 379

La Mattina, D. See A. Giambruno

Marciszewski, W. See M. Bell

Moltó, A. See R. Haydon

Mostovoy, J. and J. M. Pérez-Izquierdo Dimension filtration on loops

105

Nitzan-Hahamov, S. and A. Olevskii Sparse exponential systems: completeness with estimates

205

Olevskii, A. See S. Nitzan-Hahamov Orihuela, J. See R. Haydon Panaite, F. and M. D. Staic Generalized (anti) Yetter–Drinfeld

modules as components of a braided T-category

349

Pérez-Izquierdo, J. M. See J. Mostovoy Petrogradsky, V. M. See A. Giambruno Praeger, C. E. and C. Schneider Three types of inclusions of in-

nately transitive permutation groups into wreath products in product action

65

Quarta, L. See F. Bayart Samorodnitsky, A. See A. Barvinok Schneider, C. See C. E. Praeger

Shapira, U. Measure theoretical entropy of covers  225

Staic, M. D. See F. Panaite Ye, X. D. See W. Huang Zhang, G. See W. Huang